# Provably Efficient Reinforcement Learning with Multinomial Logit Function Approximation

### v2

@@ B.2.2 Proof of Lemma 8
This finishes the proof. ∎
B.2.3 Proof of Lemma 9
Proof.
First, notice that $\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})$ is a martingale difference sequence. Also, we have
$\quad\left|\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\right|$
$\leq\left|\left(\bar{g}_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\right|+\left|\left(g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\right|$
$\leq\lVert\bar{g}_{i,h}(\theta_{i})\rVert_{2}\lVert\theta_{i}-\theta^{*}\rVert_{2}+\lVert g_{i,h}(\theta_{i})\rVert_{2}\lVert\theta_{i}-\theta^{*}\rVert_{2}$
$\leq 4\sqrt{2}B,$
where the last inequality holds by $\lVert\bar{g}_{i,h}(\theta_{i})\rVert_{2}=\lVert\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(p_{i,h}^{s^{\prime}}(\theta)-y_{i,h}^{s^{\prime}})\phi_{i,h}^{s^{\prime}}\rVert_{2}\leq\sqrt{2}$.
We define the martingale $M_{k,h}=\sum_{i=1}^{k}\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})$, and define the conditional variance $\sigma_{i}^{2}$ as
$\sigma_{k,h}^{2}$ $=\mathbb{E}[M_{k,h}^{2}\mid\mathcal{F}_{i-1}]$
$=\sum_{i=1}^{k}\mathbb{E}_{y_{k,h}}\left[\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})^{2}\right]$
$\leq\sum_{i=1}^{k}\mathbb{E}_{y_{k,h}}\left[\left(g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})^{2}\right]$
$\leq\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*})^{2}$
$=\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}$
$\triangleq A_{k,h}.$
where the first inequality is due to the fact that $\mathbb{E}[(\xi-\mathbb{E}[\xi])^{2}]\leq\mathbb{E}[\xi^{2}]$ for any random variable $\xi$. Note that $A_{k,h}$ is a random variable, so we cannot directly apply the Bernstein inequality to $M_{k,h}$. Instead, we consider the following two cases:(i) $A_{k,h}\leq\frac{4B^{2}}{kU}$ and (ii) $A_{k,h}>\frac{4B^{2}}{kU}$.
Case (i): $A_{k,h}=\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}\leq\frac{4B^{2}}{kU}$. Then, we have
$M_{k,h}$ $=\sum_{i=1}^{k}\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})$
$=\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(p_{i,h}^{s^{\prime}}(\theta_{i})-y_{i,h}^{s^{\prime}})(\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*})$
$\leq\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lvert(\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*})\rvert$
$\leq\sqrt{kU\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}((\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*}))^{2}}$
$\leq 2B.$
where the second equality is due to the definition of $\bar{g}_{i,h}(\theta_{i})$ and $g_{i,h}(\theta_{i})$, the first inequality holds by $p_{i,h}^{s^{\prime}}(\theta_{i})-y_{i,h}^{s^{\prime}}\in[-1,1]$, the second inequality holds by the Cauchy-Schwarz inequality, and the last inequality holds by the condition of case (i).
Case (ii): $A_{k,h}=\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{j,h}}^{2}>\frac{4B^{2}}{kU}$. We have both a lower and upper bound for $A_{k,h}$, i.e., $\frac{4B^{2}}{kU}<A_{k,h}\leq 4B^{2}kU$. Then, we can use the peeling process to bound $M_{k,h}$ as follows:
$\Pr\left[M_{k,h}\geq 2\sqrt{\tau_{k,h}A_{k,h}}+\frac{8B\tau_{k,h}}{3}\right]$
=\ = $\Pr\left[M_{k,h}\geq 2\sqrt{\tau_{k,h}A_{k,h}}+\frac{8B\tau_{k,h}}{3},\frac{4B^{2}}{kU}<A_{k,h}\leq 4kUB^{2}\right]$
=\ = $\Pr\left[M_{k,h}\geq 2\sqrt{\tau_{k,h}A_{k,h}}+\frac{8B\tau_{k,h}}{3},\frac{4B^{2}}{kU}<A_{k,h}\leq 4kUB^{2},\sigma_{k,h}\leq A_{k,h}\right]$
@@ B.2.3 Proof of Lemma 9
where $m=2\log_{2}(kU)$, and the last inequality follows the Bernstein inequality for martingales.
Combining above two cases, letting $\tau=\log{\frac{mk^{2}}{\delta/H}}$ and taking the union bound over $k$ and $h\in[H]$, we have with probability at least $1-\delta$, for any $k\in[K],h\in[H]$ it holds that
$M_{k,h}$ $=\sum_{i=1}^{k}\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})$
$\leq 2\sqrt{\tau_{k,h}A_{k,h}}+\frac{8B\tau_{k,h}}{3}+4\sqrt{2}B$
$\leq 2\sqrt{\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}\log\frac{2k^{2}H\log(kU)}{\delta}}+8B\left(1+\log\frac{2k^{2}H\log(kU)}{\delta}\right)$
$=2\sqrt{\gamma_{k}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}}+8B\left(1+\gamma_{k}\right),~{}$ (23)
where $\gamma_{k}=\log\frac{2k^{2}H\log(kU)}{\delta}$.
Then, applying $uv\leq cu^{2}+v^{2}/(4c)$ for any $c,u,v>0$ with $c=2/\kappa$, we have
$\sqrt{\gamma_{k}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}}\leq\frac{2\gamma_{k}}{\kappa}+\frac{\kappa}{8}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}.$ (24)
Combining (23) and (24), we have
$\sum_{i=1}^{k}\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\leq\frac{\kappa}{4}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}+(\frac{4}{\kappa}+8B)\gamma_{k}.$
This finishes the proof. ∎
Appendix C Proof of Lemma 2
@@
The gradient of $p_{s,a}(\theta)$ is given by
By the mean value theorem, there exists $\bar{\theta}=\nu\theta_{h}^{*}+(1-\nu)\widehat{\theta}_{k,h}$ for some $\nu\in[0,1]$, such that
$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}$
=\ = $\bigg{|}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\nabla p_{s,a}^{s^{\prime}}(\bar{\theta})(\widehat{\theta}_{k,h}-\theta_{h}^{*})V(s^{\prime})\bigg{|}$
=\ = $\bigg{|}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\bar{\theta})\phi_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h}-\theta_{h}^{*})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\bar{\theta})\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\bar{\theta})\phi_{s,a}^{s^{\prime\prime}}(\widehat{\theta}_{k,h}-\theta_{h}^{*})V(s^{\prime})\bigg{|}$
=\ = $\bigg{|}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\bar{\theta})\phi_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h}-\theta_{h}^{*})\Big{(}V(s^{\prime})-\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\bar{\theta})V(s^{\prime\prime})\Big{)}\bigg{|}$
[…]
where the first inequality is by $V:\mathcal{S}\to[0,H]$, the second inequality is by the definition of $\phi_{k,h}^{s^{\prime}}$, the third inequality is by the Holder’s inequality, the last inequality is by Lemma 1. ∎
Appendix D Proof of Theorem 1
D.1 Main Proof
To prove the theorem, we first introduce the following lemma.
Lemma 10.
Suppose for all $(h,s,a)\in[H]\times\mathcal{S}\times\mathcal{A}$, it holds that
$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}\leq\Gamma_{h,s,a}.$ (26)
Define
$\widehat{Q}_{k,h}(s,a)=\bigg{[}r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})\widehat{V}_{k,h+1}(s^{\prime})+\Gamma_{h,s,a}\bigg{]}_{[0,H]}.$ (27)
Then, for any $\delta\in(0,1]$, with probability at least $1-\delta$, it holds that
$\operatorname*{Reg}(K)\leq 2\sum_{k=1}^{K}\sum_{h=1}^{H}\Gamma_{h,s_{k,h},a_{k,h}}+H\sqrt{2KH\log(2/\delta)}.$
Proof of Theorem 1.
Substituting $\Gamma_{h,s,a}=H\widehat{\beta}_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{k,h}^{-1}}$ into Lemma 10, we have
$\operatorname*{Reg}(K)$ $\leq 2H\widehat{\beta}_{k}\sum_{k=1}^{K}\sum_{h=1}^{H}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{k,h}^{-1}}+H\sqrt{2KH\log(2/\delta)}$
$\leq 2H\widehat{\beta}_{k}\sum_{h=1}^{H}\sqrt{K\sum_{k=1}^{K}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\widehat{\Sigma}_{k,h}^{-1}}^{2}}+H\sqrt{2KH\log(2/\delta)}$
$\leq 2H^{2}\widehat{\beta}_{k}\sqrt{\frac{4dK}{\kappa}\log\left(1+\frac{KU}{d\lambda}\right)}+H\sqrt{2KH\log(2/\delta)}\leq\widetilde{\mathcal{O}}({\kappa}^{-1}dH^{2}\sqrt{K}),$
where the second inequality holds by the Cauchy-Schwarz inequality and the third inequality holds by Lemma 6. This finishes the proof. ∎
D.2 Proof of Auxiliary Lemmas
In this section, we provide the proofs of the lemmas used in Appendix D.1.
First, we introduce the following lemma.
Lemma 11.
Suppose (26) and (27) in Lemma 10 holds for all $k\in[K],h\in[H]$. Then, for any $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times[H]$, it holds that
$Q_{h}^{*}(s,a)\leq\widehat{Q}_{k,h}(s,a)\leq r_{h}(s,a)+\mathbb{P}_{h}V_{k,h+1}(s,a)+2\Gamma_{h,s,a}.$
@@ D.2 Proof of Auxiliary Lemmas
First, we prove the left-hand side of the lemma. We prove this by backward induction on $h$. For the stage $h=H$, by definition, we have
$\widehat{Q}_{k,H}(s,a)=r_{H}(s,a)=Q_{H}^{*}(s,a),\quad\widehat{V}_{k,H+1}(s)=0=V_{H+1}^{*}(s).$
Suppose the statement holds for $h+1$, we show it holds for $h$. By definition, if $\widehat{Q}_{k,h}(s,a)=H$, this holds trivially. Otherwise, we have
$\widehat{Q}_{k,h}(s,a)$ $=r_{h}(s,a)+p_{s,a}(\widehat{\theta}_{k,h})\widehat{V}_{k,h+1}+\Gamma_{h,s,a}$
$\geq r_{h}(s,a)+p_{s,a}(\widehat{\theta}_{k,h})V_{k}^{*}+\Gamma_{h,s,a}\geq r_{h}(s,a)+p_{s,a}(\theta_{h}^{*})V_{k}^{*}=Q_{h}^{*}(s,a).$
Then, we prove the right-hand side of the lemma. By the definition of $\widehat{Q}_{k,h}(s,a)$, we have
$\widehat{Q}_{k,h}(s,a)=r_{h}(s,a)+p_{s,a}(\widehat{\theta}_{k,h})\widehat{V}_{k,h+1}+\Gamma_{h,s,a}\leq r_{h}(s,a)+p_{s,a}(\theta_{h}^{*})\widehat{V}_{k,h+1}+2\Gamma_{h,s,a},$
where the inequality is by (26). ∎
D.2.1 Proof of Lemma 10
Proof.
By the definition of $\operatorname*{Reg}(K)=\sum_{k=1}^{K}V_{1}^{*}(s_{k,1})-\sum_{k=1}^{K}V_{1}^{\pi_{k}}(s_{k,1})$, we have
@@ D.2.1 Proof of Lemma 10
where the first inequality is by Lemma 11 and the second is by $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s_{k,h},a)$.
@@
=\ = $r(s_{k,1},a_{k,1})+\mathbb{P}_{1}\widehat{V}_{k,2}(s_{k,1},a_{k,1})++2\Gamma_{h,s_{k,1},a_{k,1}}-r(s_{k,1},a_{k,1})-\mathbb{P}_{1}V_{2}^{\pi_{k}}(s_{k,1},a_{k,1})$
@@
Summing over $k$, we have for any $\delta\in(0,1]$, with probability at least $1-\delta$, it holds that
@@
Appendix E Proof of Lemma 3

### v3

@@ C.2 Proof of Lemma 7
$+\frac{3H}{2}\sum_{s^{\prime}}\big{(}u_{s,a}^{s^{\prime}}(\theta_{2})-u_{s,a}^{s^{\prime}}(\theta_{1})\big{)}^{2}p_{s,a}^{s^{\prime}}(u(\bar{\theta})).$ (17)
To bound the first term, by applying the AM-GM inequality, we obtain
$H\sum_{s^{\prime}}\sum_{s^{\prime\prime}\neq s^{\prime}}\big{\lvert}u_{s,a}^{s^{\prime}}(\theta_{2})-u_{s,a}^{s^{\prime}}(\theta_{1})\big{\rvert}\cdot p_{s,a}^{s^{\prime}}\big{(}u(\bar{\theta})\big{)}p_{s,a}^{s^{\prime\prime}}\big{(}u(\bar{\theta})\big{)}\cdot\big{\lvert}u_{s,a}^{s^{\prime\prime}}(\theta_{2})-u_{s,a}^{s^{\prime\prime}}(\theta_{1})\big{\rvert}$
@@
$\hskip 56.9055pt+\frac{H}{2}\sum_{s^{\prime}}\sum_{s^{\prime\prime}}\big{(}u_{s,a}^{s^{\prime\prime}}(\theta_{2})-u_{s,a}^{s^{\prime\prime}}(\theta_{1})\big{)}^{2}p_{s,a}^{s^{\prime}}\big{(}u(\bar{\theta})\big{)}p_{s,a}^{s^{\prime\prime}}\big{(}u(\bar{\theta})\big{)}$
@@
Plugging (18) into (17), we have
$\frac{H}{2}\sum_{s^{\prime}}\sum_{s^{\prime\prime}}\big{\lvert}u_{s,a}^{s^{\prime}}(\theta_{2})-u_{s,a}^{s^{\prime}}(\theta_{1})\big{\rvert}\cdot\frac{\partial^{2}F(u(\bar{\theta}))}{\partial s^{\prime}\partial s^{\prime\prime}}\cdot\big{\lvert}u_{s,a}^{s^{\prime\prime}}(\theta_{2})-u_{s,a}^{s^{\prime\prime}}(\theta_{1})\big{\rvert}$
[…]
=\ = $\frac{5H}{2}\sum_{s^{\prime}}\big{(}(\phi_{s,a}^{s^{\prime}})^{\top}(\theta_{2}-\theta_{1})\big{)}^{2}p_{s,a}^{s^{\prime}}(u(\bar{\theta}))$
[…]
where the last inequality holds by the condition $\lVert\theta_{1}-\theta_{2}\rVert_{\Lambda}\leq\beta$.
Finally, combining (16) and (19) finishes the proof. ∎
C.3 Proof of Lemma 8
@@
First, we prove the left-hand side of the lemma. We prove this by backward induction on $h$. For the stage $h=H$, by definition, we have $\widehat{Q}_{k,H}(s,a)=r_{H}(s,a)=Q_{H}^{*}(s,a),\widehat{V}_{k,H+1}(s)=0=V_{H+1}^{*}(s)$. Suppose the statement holds for $h+1$, we show it holds for $h$. By definition, if $\widehat{Q}_{k,h}(s,a)=H$, this holds trivially. Otherwise, we consider two cases:
For $\widehat{Q}_{k,h}(s,a)$ defined in (14), we have
$\widehat{Q}_{k,h}(s,a)$ $=r_{h}(s,a)+\operatorname*{arg\,max}_{\theta\in\mathcal{C}_{k,h}}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta)\widehat{V}_{k,h+1}(s^{\prime})$
$\geq r_{h}(s,a)+\operatorname*{arg\,max}_{\theta\in\mathcal{C}_{k,h}}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta)V_{k,h+1}^{*}(s^{\prime})$
$\geq r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V_{k,h+1}^{*}(s^{\prime})=Q_{h}^{*}(s,a).$
where the first inequality is by the induction hypothesis, and the second inequality is due to $\theta_{h}^{*}\in\mathcal{C}_{k,h}$.
For $\widehat{Q}_{k,h}(s,a)$ defined in (15), we have
$\widehat{Q}_{k,h}(s,a)$ $=r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}(\widehat{\theta}_{k,h})\widehat{V}_{k,h+1}(s^{\prime})+\Gamma_{h,s,a}$
$\geq r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}(\widehat{\theta}_{k,h})V_{k,h+1}^{*}(s^{\prime})+\Gamma_{h,s,a}$
$\geq r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}(\theta_{h}^{*})V_{k,h+1}^{*}(s^{\prime})=Q_{h}^{*}(s,a).$
Then, we prove the right-hand side of the lemma.
For $\widehat{Q}_{k,h}(s,a)$ defined in (14), we have
$\widehat{Q}_{k,h}(s,a)$ $=r_{h}(s,a)+\operatorname*{arg\,max}_{\theta\in\mathcal{C}_{k,h}}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta)\widehat{V}_{k,h+1}(s^{\prime})$
$\leq r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})\widehat{V}_{k,h+1}(s^{\prime})+\Gamma_{k,h,s,a}$
$\leq r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})\widehat{V}_{k,h+1}(s^{\prime})+2\Gamma_{h,s,a},$
where the inequality is by (13).
For $\widehat{Q}_{k,h}(s,a)$ defined in (15), we have
$\widehat{Q}_{k,h}(s,a)$ $=r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})\widehat{V}_{k,h+1}(s^{\prime})+\Gamma_{k,h,s,a}$
$\leq r_{h}(s,a)+\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})\widehat{V}_{k,h+1}(s^{\prime})+2\Gamma_{k,h,s,a},$
where the inequality is by (13). ∎
C.4 Proof of Lemma 9
@@
By the definition that $\operatorname*{Reg}(K)=\sum_{k=1}^{K}V_{1}^{*}(s_{k,1})-\sum_{k=1}^{K}V_{1}^{\pi_{k}}(s_{k,1})$, we have for any $\delta\in(0,1]$, with probability at least $1-\delta$, it holds that
@@
where the first inequality is by Lemma 8 and $\theta_{h}^{*}\in\widehat{\mathcal{C}}_{k,h}$ with probability at least $1-\delta$, and the second inequality is by action selection $a_{k,h}=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widehat{Q}_{k,h}(s_{k,h},a)$.
@@
=\ = $r(s_{k,1},a_{k,1})+\mathbb{P}_{1}\widehat{V}_{k,2}(s_{k,1},a_{k,1})+2\Gamma_{h,s_{k,1},a_{k,1}}-r(s_{k,1},a_{k,1})-\mathbb{P}_{1}V_{2}^{\pi_{k}}(s_{k,1},a_{k,1})$
@@
Summing over $k$, we have
@@
Appendix D Omitted Proofs for Section 3
D.1 Proof of Claim 1